Divide P(x) by D(x) using synthetic division. Write P(x) in the form Q(x)D(x) + R(x).


  1. P(x) = x4 + x3 - 3x2 - 5x - 2; D(x) = x - 2
  2. P(x) = x3 - 9x - 12; D(x) = x + 3
  3. P(x) = -2x3 - 5x4 + 2x5 - 3x + 1; D(x) = 2x - 1
  4. P(x) = 4x4 - 3x2 + x - 17; D(x) = 2x + 3







1
Expert's answer
2021-01-27T15:26:43-0500

1)


"\frac{ ( x^4 + x^3 - 3x^2 - 5x - 2)}{( x - 2)}=x^3 + 3x^2 +3x +1"

2)


"\frac{ ( x^3 - 9x - 12)}{( x+3)}=x^2 -3x -\frac{12}{( x+3)}"

3)


"\frac{ ( -2x^3 - 5x^4 + 2x^5 - 3x + 1)}{( 2x-1)}=x^4 -2x^3-2x^2 -x-2\\-\frac{ 1}{( 2x-1)}"

4)


"\frac{ ( 4x^4 -3x^2+x-17)}{( 2x+3)}=2x^3-3x^2+3x-4\\-\frac{ 5}{( 2x+3)}"